The revenue for selling x coffee makers is given by R = x(43 – 0.2x) dollars. The cost of each coffee maker is $19. How many coffee makers must be sold to achieve a profit of at least $675.00? (Hint: Profit = Revenue - Cost)
Answer by stanbon(75887)

The revenue for selling x coffee makers is given by R = x(43 – 0.2x) dollars. The cost of each coffee maker is $19. How many coffee makers must be sold to achieve a profit of at least $675.00? (Hint: Profit = Revenue - Cost)
-------------
INEQUALITY:
x(43-0.2x)-19x >= 675
43x-0.2x^2-19x >= 675
-0.2x^2+24x-675 >=0
I graphed to find the solution:
Ans: 45 <= x <= 75
======================
Cheers,
Stan H.